OpenStudy (anonymous):

Rewrite parabola equations in standard form: 1. x^2+14x+44y+313=0 2. 3y^2+6y+108x-969=0 3. x^2-12x-48y-373=0

OpenStudy (jdoe0001):

solve each for "y"

OpenStudy (jdoe0001):

standard form => \(\bf y = ax^2\pm bx\pm c\)

OpenStudy (anonymous):

I'm so confused... Can you please show me how to do the first one?

OpenStudy (anonymous):

My teacher says I need to start by finding the vertex first...

OpenStudy (saifoo.khan):

There are many ways to find the equation. You may use whichever you like.

OpenStudy (anonymous):

Can you please help me and show me how to do the first one by finding the vertex?

OpenStudy (jdoe0001):

hmm the "vertex form" and the "standard form" are no exactly the same.... -> http://www.mathwarehouse.com/geometry/parabola/images/standard-vertex-forms-thumb.png maybe you're just asked to get the vertex form?

OpenStudy (anonymous):

In order to find the standard form she says I must find the vertex and focus first, but answers are in fact supposed to be in standard form

OpenStudy (jdoe0001):

or maybe he/she meant \(\Large \bf (x-h)^2=4p(y-k) \quad ?\)

OpenStudy (jdoe0001):

ahhh the vertex and focus. point... then yes.... that will be \(\large \bf (x-h)^2=4p(y-k)\)

OpenStudy (anonymous):

Yes thats it! How do I do it? haha

OpenStudy (jdoe0001):

you know what a "perfect square trinomial" is, right?

OpenStudy (anonymous):

I don't think so...

OpenStudy (jdoe0001):

hmm lemme do a quick rundown on that... a perfect square trinomial is a TRInomial thus 3 terms it comes from multiplying 2 EQUAL binomials so \(\Large \bf (a\pm b)(a\pm b)\to (a\pm b)^2=a^2\pm 2ab+b^2\)

OpenStudy (jdoe0001):

see the TRInomial?

OpenStudy (anonymous):

Yes, that makes sense

OpenStudy (jdoe0001):

so the trinomial rule is pretty simple if you look at the two terms in the binomial so \(\Large \bf ({\color{blue}{ a}}\pm {\color{brown}{ b}})^2={\color{blue}{ a}}^2\pm 2\cdot {\color{blue}{ a}}{\color{brown}{ b}}+{\color{brown}{ b}}^2\) notice the middle term is really just 2 times both terms and notice the left and right sides of the trinomial is just each term squared

OpenStudy (jdoe0001):

so with that in mind.... let us grab the 1st equation

OpenStudy (anonymous):

Ok perfect

OpenStudy (jdoe0001):

\(\bf x^2+14x+44y+313=0\implies x^2+14x=-44y-313 \\ \quad \\ (x^2+14x+{\color{red}{ \square ? }}^2)=-44y-313\) any ideas what our missing fellow there is? wel really need a number there to "complete the square" so we need a number there, that multiplied times the 1st term and 2, will give us the middle term

OpenStudy (anonymous):

49?

OpenStudy (jdoe0001):

49? let's try that

OpenStudy (anonymous):

Oh wait, so it would have to be a number squared... um maybe 7?

OpenStudy (jdoe0001):

\(\large x^2+14x+{\color{red}{ 49 }}^2\implies \begin{cases} middle\ term\\ \hline\\ 2\cdot x\cdot 7\implies 14x \end{cases}\ yeap, is 49, or \(7^2\) now, keep in mind that all we're doing is really "borrowing from our good friend, Mr Zero, 0" so if we ADD \(\7^2\) we also have to SUBTRACT \(\7^2\) so \(\bf x^2+14x+44y+313=0\implies x^2+14x=-44y-313 \\ \quad \\ (x^2+14x+{\color{red}{ 7 }}^2)-{\color{red}{ 7 }}^2=-44y-313 \\ \quad \\ (x+7)^2-49=-44y-313\implies (x+7)^2=-44y-313+49 \\ \quad \\ (x+7)^2=-44y-264\implies (x+7)^2=-44(y-6)\)

OpenStudy (jdoe0001):

hmm anyhow .. tis a bit broken but \(\bf \large x^2+14x+{\color{red}{ 49 }}^2\implies \begin{cases} middle\ term\\ \hline\\ 2\cdot x\cdot 7\implies 14x \end{cases}\) yeap, is 49, or \(7^2\) so we add \(7^2\) and we subtract \(7^2\)

OpenStudy (anonymous):

So, (x+7)^2=-44(y-6) How would I get the final answer from here, or where does the 4 come in from (x-h)^2=4p(y-k)?

OpenStudy (jdoe0001):

hmmm should be +6 actually since the negative is outside... so \(\bf x^2+14x+44y+313=0\implies x^2+14x=-44y-313 \\ \quad \\ (x^2+14x+{\color{red}{ 7 }}^2)-{\color{red}{ 7 }}^2=-44y-313 \\ \quad \\ (x+7)^2-49=-44y-313\implies (x+7)^2=-44y-313+49 \\ \quad \\ (x+7)^2=-44y-264\implies (x+7)^2=-44(y+6)\)

OpenStudy (jdoe0001):

hmm ohh yeah... the 4... well the form is really \(\bf (x+7)^2=-44(y+6)\) the "4p" just happen to be equal to -44

OpenStudy (anonymous):

Ok! Thank you so much! It makes sense now, but I'm a little confused on the second one now with the 3y^2 and etc

OpenStudy (jdoe0001):

the 2nd one is the same notice the squared variable however is the "y" NOT the "x" that just means is a "horizontal" parabola